# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ 2 Outline of Converse Proofs
First term. The thresholds in (4) and (11) consist of two terms. The converse for the first term maximized over $\nu$ is known even for arbitrary test designs via a simple channel capacity argument [33, 38], and for arbitrary $\nu$, the converse for Bernoulli designs is also known via [39]. We required non-minor additional technical effort to handle the near-constant weight design with arbitrary $\nu$, [^3] A simpler approach based on Fano’s inequality (e.g., see [33]) could also be used, but would only give the weaker statement $\mathbb{P}({\rm err})\not\to 0$ rather than $\mathbb{P}({\rm err})\to 1$. but the intuition behind the bound is simple: With parameter $\nu$, each test has probability roughly $e^{-\nu}$ of containing no defectives, and thus a probability roughly $e^{-\nu}\star\rho$ of being positive. A standard information-theoretic argument shows that each test can only reveal $H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho)$ bits of information. On the other hand, with (${p\choose k}$ possible defective sets, we require roughly $\log{p\choose k}=\big{(}k\log_{2}\frac{p}{k}\big{)}(1+o(1))$ bits to identify the correct one. Comparing these quantities leads to the first term in (11). This part of the analysis will be detailed in Appendix I.
Discussion of second term. The second terms in (10) and (17) are dominant at high values of $\theta$. Terms of this kind (albeit much simpler ones) are widely known in the noiseless setting [3, 4, 29, 17], and are based on the idea of that if a defective item is masked (i.e., every test it is in also contains at least one other defective), then even an optimal decoder will be unable to identify it. Intuitively, this is because the tests results are unchanged when that item is changed to be non-defective.
In the noisy setting, complete masking is no longer the dominant error event, but we can use a similar idea: Informally, an optimal decoder will fail when the following events hold simultaneously (see Corollary 1 in Appendix A for the formal version):
@@
The idea is that when both of these occur with suitable notions of “few” and “many”, the set $(S\setminus\{j\})\cup\{j^{\prime}\}$ will be favored to $S$ itself. We proceed with further details towards making this intuition rigorous, and we will see that the formal version of conditions (i) and (ii) above lead to the terms $f_{1}$ and $f_{2}$ in the final results. This part of the analysis will be detailed in Appendix C (Bernoulli design) and Appendix E (near-constant weight design).
Maximum-likelihood decoding. Under a uniform prior on $S$, it is well known that the optimal decoding rule is maximum-likelihood decoding. Under our symmetric noise model, this is equivalent to choosing an estimate $\hat{S}$ such that as many tests as possible match what the noiseless outcome would be. Thus, referring back to the above outline, we seek to find $j\in S$ and $j^{\prime}\notin S$ such that $(S\setminus\{j\})\cup\{j^{\prime}\}$ has more such tests than $S$ itself.
Identifying a suitable defective item. Suppose without loss of generality that $S=\{1,\dotsc,k\}$. We fix parameters $C>0$ and $\zeta\in(0,1)$ to be optimized at the end of our analysis (thus appearing as optimization parameters in our theorems), and consider the following “bad event”: [^4] This parametrization of $C$ turns out to be more notationally convenient. A given defective item has $\frac{Cn\nu e^{-\nu}}{k}$ tests in which it is the only defective, and a fraction $\zeta$ of those tests are flipped. We show that if $n$ is below the stated threshold depending on $f_{1}$, then some $j\in\{1,\dotsc,k\}$ satisfies this with high probability.
@@
  • Interpret the placements of items into tests as edges in a bipartite graph (or more precisely a multi-graph, since the same item can be placed in the same test twice).
  • Establish that among the $k\Delta$ edges connecting defective items to tests, roughly $e^{-\nu}k\Delta$ of them are to tests containing exactly one defective item.
@@
Consider any fixed $j^{\prime}\in\{k+1,\dotsc,p\}$. When comparing the likelihoods of $(S\setminus\{j\})\cup\{j^{\prime}\}$ and $S$, there are only two types of tests that ultimately matter:
@@
This is because all other tests contribute the same amount to the likelihood even after removing $j$ and adding $j^{\prime}$ (due to the “OR” operation in (1)). We are interested in counting how many tests of the above kind $j^{\prime}$ is placed in, and moreover, how many of those tests are negative vs. positive (both are possible due to the noise). This comes down to the analysis of various binomial distributions, though it is complicated by delicate dependencies. We show that tests of the second kind above are dominant, whereas in this part of the analysis, the tests of the first kind above only contribute to lower-order asymptotic terms.
@@ 3 Outline of Achievability Proofs
In this section, we outline the proofs of our achievability results (i.e., upper bounds on the required number of tests), which come with several additional challenges compared to the converse results.
@@ 3.1 Information-Theoretic Tools
Let $\mathbf{X}\in\{0,1\}^{n\times p}$ denote the test matrix, $\mathbf{Y}\in\{0,1\}^{n}$ the test results, and $S\subseteq[p]$ the defective set. Since both test designs are symmetric with respect to re-ordering items, we may focus on a specific choice of $S=s$ in our achievability analysis, say $s=[k]:=\{1,\dotsc,k\}$. Let $\mathbf{X}_{s}$ denote the resulting $n\times k$ sub-matrix of $\mathbf{X}$, and similarly when $s$ is replaced by any other subset of $[p]:=\{1,\dotsc,p\}$.
An error occurs when some $\bar{s}\neq s$ is favored by the decoder (to be defined below), and the contribution to the error probability can differ significantly depending on the amount of overlap between $\bar{s}$ and $s$. For instance, there are only relatively few sets $\bar{s}$ with $|s\setminus\bar{s}|=1$, but the individual probability of favoring such $\bar{s}$ is relatively high due to the low overlap. To capture this, we consider partitioning $s$ into $(s_{\mathrm{dif}},s_{\mathrm{eq}})$ with $s_{\mathrm{dif}}\neq\emptyset$, and we denote $\ell=|s_{\mathrm{dif}}|\in\{1,\dotsc,k\}$ (so that $|s_{\mathrm{eq}}|=k-\ell$). Intuitively, $s_{\mathrm{eq}}$ represents where an incorrect estimate overlaps with $s$, and $s_{\mathrm{dif}}$ represents the differing part.
@@
$\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}):=\log\frac{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{dif}}},\mathbf{X}_{s_{\mathrm{eq}}})}{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})},$ (18)
which is the log-likelihood ratio of $\mathbf{Y}$ given the full test sub-matrix $\mathbf{X}_{s}$ vs. the smaller sub-matrix $\mathbf{X}_{s_{\mathrm{eq}}}$ alone. Following the extensive literature on similar techniques for channel coding [25], we refer to this quantity as the information density. Notice that the average of (18) with respect to $(\mathbf{X},\mathbf{Y})$ is the following conditional mutual information:
$I_{\ell}^{n}:=I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}),$ (19)
@@
$\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})\geq\gamma_{|s_{\mathrm{dif}}|},\quad\forall(s_{\mathrm{dif}},s_{\mathrm{eq}})\text{ such that }|s_{\mathrm{dif}}|\neq 0.$ (20)
If multiple such $s$ exist, or if none exist, then an error is declared. This decoder is inspired by analogous thresholding techniques from the channel coding literature [23, 25], with the rough idea being that the numerator in (18) (i.e., the overall likelihood) tends to be much higher than the denominator for the correct set $s$, whereas for an incorrect $s$ that has overlap $s_{\mathrm{eq}}$ with the correct one, such behavior is highly unlikely.
Limitation of existing approach. While the analysis of the decoder (20) in [39] leads to optimal thresholds in the noiseless setting, we found that even a sharpened analysis under their framework leads to a suboptimal result in the noisy setting. In more detail, their analysis leads to three terms: (i) a mutual information based term for $\ell=k$; (ii) a mutual information based term for $\ell=1$; and (iii) a term capturing the concentration behavior of the information density. We found that a tight concentration analysis can lead to improvements in the third of these. [^5] Namely, in the denominator of the final result we get $\nu e^{-\nu}(1-e^{-D(1/2\|\rho)})$ for the Bernoulli design, and $\nu^{2}-\nu\log(e^{\nu}-1+e^{-D(1/2\|\rho)})$ for the near-constant weight design. We state these without proof because they would significantly lengthen the paper but still give a suboptimal final result. However, there still exists a parameter (called $\delta_{2}$ in [39]) that trades off the second and third terms and leaves the result significantly suboptimal, and we were unable to identify any promising route to avoiding this limitation when using the framework of [39].
@@ 3.2 A Hybrid Decoding Rule
    $\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s})>\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s^{\prime}}),\quad\forall s^{\prime}\text{ such that }1\leq|s\setminus s^{\prime}|\leq\frac{k}{\log k},$ (21)
@@
    $\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})\geq\gamma_{|s_{\mathrm{dif}}|},\quad\forall(s_{\mathrm{dif}},s_{\mathrm{eq}})\text{ such that }|s_{\mathrm{dif}}|>\frac{k}{\log k},$ (22)
@@ 3.3 Analysis of the Low Overlap (High $\ell$) Regime
For the Bernoulli design, the analysis of the threshold decoder (22) for $\ell>\frac{k}{\log k}$ will be taken directly from [39], and no change is needed. However, most of the analysis in [39] relied heavily on the test matrix having independent rows, so it is not applicable to the near-constant weight design. Thus, we need to analyze the information density and mutual information to fill in these gaps, with the main steps being as follows:
@@ 3.4 Analysis of the High Overlap (Low $\ell$) Regime
For the $-\ell$ regime, we follow a similar structure to the converse bound, but we now need to consider all $\ell=1,\dotsc,\frac{k}{\log k}$ instead of only $\ell=1$. Specifically, we again consider (optimal) maximum-likelihood decoding, and accordingly, we seek to show that $S$ is preferred to any other set of the form $(S\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$ with $\mathcal{J}\subset S$ and $\mathcal{J}^{\prime}\subset\{1,\dotsc,p\}\setminus S$ both having cardinality $1\leq\ell\leq\frac{k}{\log k}$. Recall that “being preferred” is equivalent to having more tests that match what the noiseless outcome would have been.
@@
For the $(\ell,C,\zeta)$ triplets for which we cannot guarantee $k_{\ell,C,\zeta}=0$, we need to consider the placements of non-defectives. We again follow a similar argument to the converse, but with more general $\ell\geq 1$ and all relevant $(C,\zeta)$ pairs being considered simultaneously. In particular, moving from $\ell=1$ to $\ell>1$ adds significant additional technical challenges, but conceptually we still follow similar steps as in the converse analysis. The term corresponding to $\ell=1$ turns out to be the dominant one, and this part of the analysis leads to the term $f_{2}$ in the final bound.
@@ 4 Conclusion
We have derived exact asymptotic thresholds on the number of tests required for probabilistic group testing under binary symmetry noise, under both the Bernoulli design and the near-constant weight design. We believe that our results bring the degree of understanding of this noisy group testing model significantly closer to the noiseless setting, but two major open problems remain:
  • It is currently unclear whether the exact threshold for the near-constant weight design (with optimized $\nu$) is also asymptotically optimal among all non-adaptive designs.
  • The upper bounds were only derived using computationally intractable information-theoretic decoding rules, and establishing the same thresholds using polynomial-time algorithms (possibly with the test design slightly modified, e.g., via spatial coupling techniques [17]) would be of significant interest.
Noiseless counterparts for both of these points were established in [17], so their techniques may serve as the natural starting point for studying the same questions in the noisy setting.

### 23 Oct 2024 v2

@@ 2 Outline of Converse Proofs
Discussion of the first term. The thresholds in (1.4) and (1.11) consist of two terms. The converse for the first term maximized over $\nu$ is known even for arbitrary test designs via a simple channel capacity argument [33, 38], and for arbitrary $\nu$, the converse for Bernoulli designs is also known via [39]. We required non-minor additional technical effort to handle the near-constant weight design with arbitrary $\nu$, [^4] A simpler approach based on Fano’s inequality (e.g., see [33]) could also be used, but would only give the weaker statement $\mathbb{P}({\rm err})\not\to 0$ rather than $\mathbb{P}({\rm err})\to 1$. but the intuition behind the bound is simple: With parameter $\nu$, each test has probability roughly $e^{-\nu}$ of containing no defectives, and thus a probability roughly $e^{-\nu}\star\rho$ of being positive. A standard information-theoretic argument then shows that each test can only reveal $H_{2}(e^{-\nu}\star\rho)-H_{2}(\rho)$ bits of information. On the other hand, with (${p\choose k}$ possible defective sets, we require roughly $\log{p\choose k}=\big{(}k\log_{2}\frac{p}{k}\big{)}(1+o(1))$ bits to identify the correct one. Comparing these quantities leads to the first term in (1.11). This part of the analysis will be detailed in Appendix I.
Discussion of the second term. The second terms in (1.10) and (1.16) are dominant at high values of $\theta$. Terms of this kind (albeit much simpler ones) are widely known in the noiseless setting [3, 4, 29, 16], and are based on the idea of that if a defective item is masked (i.e., every test it is in also contains at least one other defective), then even an optimal decoder will be unable to identify it. Intuitively, this is because the tests results are unchanged when that item is changed to be non-defective.
In the noisy setting, complete masking is no longer the dominant error event, but we can use a similar idea. Informally, an optimal decoder will fail when the following events hold simultaneously (see Corollary B.1 in Appendix B for the formal version):
@@
The idea is that when both of these occur with suitable notions of “few” and “many”, the set $(S\setminus\{j\})\cup\{j^{\prime}\}$ will be favored to $S$ itself. We proceed with further details towards making this intuition rigorous, and we will see that the formal version of conditions (i) and (ii) above lead to the terms $f_{1}$ and $f_{2}$ in the final results. This part of the analysis will be detailed in Appendix C (Bernoulli design) and Appendix E (near-constant weight design), and in the following, we provide a more detailed outline of some of the main ideas and analysis techniques that we use.
Maximum-likelihood decoding. Under a uniform prior on the defective set $S$, it is well known that the optimal decoding rule is maximum-likelihood decoding. Under our symmetric noise model, this is equivalent to choosing an estimate $\hat{S}$ such that as many tests as possible match what the noiseless outcome would be. Thus, referring back to the above outline, we seek to find $j\in S$ and $j^{\prime}\notin S$ such that $(S\setminus\{j\})\cup\{j^{\prime}\}$ has more such tests than $S$ itself.
Identifying a suitable defective item. Suppose without loss of generality that $S=\{1,\dotsc,k\}$. We fix parameters $C>0$ and $\zeta\in(0,1)$ to be optimized at the end of our analysis (thus appearing as optimization parameters in our theorems), and consider the following “bad event”: [^5] The parametrization $\frac{Cn\nu e^{-\nu}}{k}$ turns out to be more notationally convenient than $\frac{C^{\prime}n}{k}$, so we adopt the former. A given defective item has $\frac{Cn\nu e^{-\nu}}{k}$ tests in which it is the only defective, and a fraction $\zeta$ of those tests are flipped. We show that if $n$ is below the stated threshold depending on $f_{1}$, then some $j\in\{1,\dotsc,k\}$ satisfies this with high probability.
@@
  • Interpret the placements of items into tests as edges in a bipartite graph (or more precisely a multi-graph, since a given item can be placed in the same test more than once).
  • Establish that among the $k\Delta$ edges connecting defective items to tests, roughly $e^{-\nu}k\Delta$ of them are connected to tests containing exactly one defective item.
@@
Consider any fixed non-defective $j^{\prime}\in\{k+1,\dotsc,p\}$. When comparing the likelihoods of $(S\setminus\{j\})\cup\{j^{\prime}\}$ and $S$, there are only two types of tests that ultimately matter:
@@
This is because all other tests contribute the same amount to the likelihood even after removing $j$ and adding $j^{\prime}$ (due to the “OR” operation in (1.1)). We are interested in counting how many tests of the above kind $j^{\prime}$ is placed in, and moreover, how many of those tests are negative vs. positive (both are possible due to the noise). We note that the parameter $d$ (appearing in $f_{2}$) is related to the number of such tests that are positive. This part of the analysis comes down to the behavior of various binomial distributions, though it is complicated by delicate dependencies. We show that tests of the second kind above are dominant, whereas in this part of the analysis, the tests of the first kind above only contribute to lower-order asymptotic terms.
@@ 3 Outline of Achievability Proofs
In this section, we outline the proofs of our achievability results (i.e., upper bounds on the required number of tests), which come with additional challenges compared to the converse results.
@@ 3.1 Information-Theoretic Tools
Let $\mathbf{X}\in\{0,1\}^{n\times p}$ denote the test matrix, $\mathbf{Y}\in\{0,1\}^{n}$ the test results, and $S\subseteq[p]$ the defective set. Since both test designs are symmetric with respect to re-ordering items, we may focus on a specific choice of $S=s$ in our achievability analysis, say $s=[k]:=\{1,\dotsc,k\}$. Let $\mathbf{X}_{s}$ denote the resulting $n\times k$ sub-matrix of $\mathbf{X}$ obtained by taking the columns indexed by $s$, and similarly when $s$ is replaced by any other subset of $[p]:=\{1,\dotsc,p\}$.
An error occurs when some $\bar{s}\neq s$ is favored by the decoder (to be defined below), and the contribution to the error probability can differ significantly depending on the amount of overlap between $\bar{s}$ and $s$. For instance, there are only relatively few sets $\bar{s}$ with $|s\setminus\bar{s}|=1$, but the individual probability of favoring such $\bar{s}$ is relatively high due to the low overlap. To capture this, we consider partitioning $s$ into $(s_{\mathrm{dif}},s_{\mathrm{eq}})$ with $s_{\mathrm{dif}}\neq\emptyset$, and we denote $\ell=|s_{\mathrm{dif}}|\in\{1,\dotsc,k\}$ (so that $|s_{\mathrm{eq}}|=k-\ell$). Intuitively, $s_{\mathrm{eq}}$ represents where an incorrect estimate overlaps with $s$, and $s_{\mathrm{dif}}$ represents the differing part, and $\ell=|s_{\mathrm{dif}}|$ represents the amount of deviation.
@@
$\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}):=\log\frac{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{dif}}},\mathbf{X}_{s_{\mathrm{eq}}})}{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})},$ (3.1)
which is the log-likelihood ratio of $\mathbf{Y}$ given the full test sub-matrix $\mathbf{X}_{s}$ vs. the smaller sub-matrix $\mathbf{X}_{s_{\mathrm{eq}}}$ alone. Following the extensive literature on similar techniques for channel coding [25], we refer to this quantity as the information density. Notice that the average of (3.1) with respect to $(\mathbf{X},\mathbf{Y})$ is the following conditional mutual information:
$I_{\ell}^{n}:=I(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}}),$ (3.2)
@@
$\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})\geq\gamma_{|s_{\mathrm{dif}}|},\quad\forall(s_{\mathrm{dif}},s_{\mathrm{eq}})\text{ such that }|s_{\mathrm{dif}}|\neq 0.$ (3.3)
If multiple such $s$ exist, or if none exist, then an error is declared. This decoder is inspired by analogous thresholding techniques from the channel coding literature [23, 25], with the rough idea being that the numerator in (3.1) (i.e., the overall likelihood) tends to be much higher than the denominator for the correct set $s$, whereas for an incorrect $s$ that has overlap $s_{\mathrm{eq}}$ with the correct one, such behavior is highly unlikely.
Limitation of existing approach. While the analysis of the decoder (3.3) in [39] leads to optimal thresholds in the noiseless setting, we found that even a sharpened analysis under their framework leads to a suboptimal result in the noisy setting. In more detail, their analysis leads to three terms: (i) a mutual information based term for $\ell=k$; (ii) a mutual information based term for $\ell=1$; and (iii) a term capturing the concentration behavior of the information density. We found that a tight concentration analysis can lead to improvements in the third of these. [^6] Namely, in the denominator of the final result we get $\nu e^{-\nu}(1-e^{-D(1/2\|\rho)})$ for the Bernoulli design, and $\nu^{2}-\nu\log(e^{\nu}-1+e^{-D(1/2\|\rho)})$ for the near-constant weight design. We state these without proof because they would significantly lengthen the paper but still give a suboptimal final result. However, there still exists a parameter (called $\delta_{2}$ in [39]) that trades off the second and third terms and leaves the result significantly suboptimal, and we were unable to identify any promising route to avoiding this limitation when using the framework of [39].
@@ 3.2 A Hybrid Decoding Rule
    $\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s})>\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s^{\prime}}),\quad\forall s^{\prime}\text{ such that }1\leq|s\setminus s^{\prime}|\leq\frac{k}{\log k},$ (3.4)
@@
    $\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})\geq\gamma_{|s_{\mathrm{dif}}|},\quad\forall(s_{\mathrm{dif}},s_{\mathrm{eq}})\text{ such that }|s_{\mathrm{dif}}|>\frac{k}{\log k},$ (3.5)
@@ 3.3 Analysis of the Low Overlap (High $\ell$) Regime
For the Bernoulli design, the analysis of the threshold decoder (3.5) for $\ell>\frac{k}{\log k}$ will be taken directly from [39], and no change is needed. However, most of the analysis in [39] relied heavily on the test matrix having independent rows, so it is not applicable to the near-constant weight design. Thus, we need to analyze the information density and mutual information to fill in these gaps, with the main steps being as follows:
@@ 3.4 Analysis of the High Overlap (Low $\ell$) Regime
For the $-\ell$ regime, we follow a similar structure to the converse bound, but we now need to consider all $\ell=1,\dotsc,\frac{k}{\log k}$ instead of only $\ell=1$. Specifically, we again consider (optimal) maximum-likelihood decoding, and accordingly, we seek to show that $S$ is preferred to any other set of the form $(S\setminus\mathcal{J})\cup\mathcal{J}^{\prime}$ with $\mathcal{J}\subset S$ and $\mathcal{J}^{\prime}\subset\{1,\dotsc,p\}\setminus S$ both having cardinality $\ell$ for some $\ell\in\big{\{}1,\dotsc,\frac{k}{\log k}\big{\}}$. Recall that “being preferred” is equivalent to having more tests that match what the noiseless outcome would have been.
@@
For the $(\ell,C,\zeta)$ triplets for which we cannot guarantee $k_{\ell,C,\zeta}=0$, we need to consider the placements of non-defectives. We again follow a similar argument to the converse, but with more general $\ell\geq 1$ and all relevant $(C,\zeta)$ pairs being considered simultaneously, with the help of high-probability events and union bounds. Moving from $\ell=1$ to $\ell>1$ adds particular technical challenges, but conceptually we still follow similar steps as in the converse analysis, and this part of the analysis leads to the term $f_{2}$ in the final bound.
@@ 4 Conclusion
We have derived exact asymptotic thresholds on the number of tests required for probabilistic group testing under binary symmetry noise, under both the Bernoulli design and the near-constant weight design. Our results bring the degree of understanding of this noisy group testing model significantly closer to the noiseless setting. With a spatial coupling design concurrently being shown to attain the same threshold as the near-constant weight design in a computationally efficient manner [18], we believe that the main remaining open problem is as follows: Is this threshold also asymptotically optimal among the entire class of non-adaptive designs? The analogous question is indeed true in the noiseless case [16], but answering it in the noisy case may be significantly more challenging. While an affirmative answer may be expected based on the noiseless setting, a negative answer also seems highly plausible given that different error events are dominant here; in fact, we have already observed notable differences to the noiseless setting, such as the design parameter $\nu=\log 2$ not always being optimal.